Flamo:

Find the distance between the points (11, 4) and (10, 5).

Shadow:

Use the distance formula \[D = \sqrt (x _{2-} x _{1})^2 + (y _{2} - y _{1})^2\]

Flamo:

Ye. So

Shadow:

Input the values in then solve

Flamo:

\[d = \sqrt{(5-4)}^2 + (10-11)^2\]

Flamo:

Right?

Flamo:

But the Square root goes over the whole Exqation.

Flamo:

@Hero

Flamo:

is this right?

Flamo:

\[d = \sqrt{1^2 + (-1)^2}\]

Flamo:

@Shadow Is this correct so far?

Flamo:

\[d = \sqrt{1 + -1}\]

Flamo:

\[d = \sqrt{0}\]

Shadow:

5 is x2, not y2

Shadow:

Coordinates go like this (x, y)

Shadow:

Try just type out what order you think they should be in.

Flamo:

o, Okay

Shadow:

What happened @Flamo

Flamo:

?

Flamo:

I found the problem out, (-1)^2 = 1 So its \[d = \sqrt{1 + 1}\]

Flamo:

So The answer is \[d = \sqrt{2} \]

Shadow:

Mhm, your setup was fine too. You just put the y set before the x set.

Flamo:

Yeah.

Flamo:

Thanks!

Shadow:

Mhm